Shift Summary for eve, Nov. 29, 1999

Ricardo_Barbosa (rbarbosa@fnal.gov)
Tue, 30 Nov 1999 00:27:43 -0600

Shift Summary for eve shift, 29/Nov/99

On Shift: R. J. Tesarek, R. F. Barbosa

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 9 min

Accounting of lost time: Problems with Tevatron and controled access

========================================================================
Runs ended this shift: several, but inconsequential

Tapes brought to FCC this shift:

========================================================================
Accesses during shift: Valeri, Sasha L., RickT and Ricardo
to work on VV' latches.

========================================================================
Overview of shift: waiting for beam, debugging VV' latches, and chasing
gremlins... lots of gremlins... (no adition to
Angela's comment)

========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
15:30 Shift started. No beam. And everyone is working in VV' latch

18:08 Call MCR and request to come up in muon mode

18:15 Low voltage error (S7-RR71-CAM1). Not a real problem.

20:05 Input cables for VV' latches are unplugged

20:07 Initialize Run 15089. But no beam...

21:19 Stop muon run 15089.

21:30 No beam!!! Calling Sasha L. to work with VV'.

21:33 Start cosmic run 15090. Meanwhile, people work in VV'.

22:43 MCR phones to tell that they can delivery beam.
Stop cosmic run 15090.
Initialize muon run 15091

23:00 MCR confuse about beam stop crash switch status.
Crash switch pulled out to verify problem. We find
the beam stops out and phone MCR but not before
hadrons arrived.
Stop muon run 15091.

23:02 Start muon run 15092 (with real muons, finally).

23:11 Stop muon run 15092.

23:16 Start pedestal run 15093

23:30 Stop pedestal run 15093, Initialize new data run 15094.

=======================================================================
Follow up Notes on Previous Shift

- do muon run.
- don't worry about the Ar-Eth flowmeters being low unless they drop
below
50.

========================================================================
Items to Follow up on Next Shift

If more than 1 hour downtime, call RickT about VV' latches.

-- 
+---------------------------------------+
|					|
|	Ricardo Franco Barbosa		|
|		  diz			|
|		Ate' Logo!		|
|					|
+---------------------------------------+